# Coalescing hardcore-boson condensate states with nonzero momentum

### v2

@@ 1 Introduction
On the other hand, recent years have seen a growing interest in non-Hermitian descriptions of condensed-matter systems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] It has been shown that the interplay between non-Hermiticity and interaction can give rise to exotic quantum many-body effect, ranging from non-Hermitian extensions of Kondo effect [15, 27], many-body localization [21], Fermi surface in coordinate space [28], to fermionic superfluidity [18, 29]. The cooperation between the non-Hermiticity and interaction may lead to rich quantum phases due to the peculiarity of non-Hermitian system.
@@ 2 Model and condensate states
$H$ $=$ $\sum\limits_{\alpha=1}^{3}J_{\alpha}\sum_{\mathbf{r}}\frac{1}{2}\hat{a}_{\mathbf{\ r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{e}_{\alpha}}+\mathrm{H.c.}+\sum\limits_{\alpha=1}^{3}V_{\alpha}\sum_{\mathbf{r}}\hat{n}_{\mathbf{r}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}$ (1)
$+\sum\limits_{\alpha=1}^{3}\sum_{\mathbf{r}}\left(\mu_{\alpha}\hat{n}_{\mathbf{r}}\delta_{1,m_{\alpha}}+\mu_{\alpha}^{\ast}\hat{n}_{\mathbf{r}}\delta_{N_{\alpha},m_{\alpha}}\right),$
@@
$h_{\mathbf{r}}^{\alpha}=J_{\alpha}[\frac{1}{2}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{e}_{\alpha}}+\mathrm{H.c.}+\cos(\mathbf{q\cdot e}_{\alpha})(\hat{n}_{\mathbf{r}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}-\hat{n}_{\mathbf{r}}-\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}})$
$+\frac{1}{2}\left(e^{i\mathbf{q\cdot e}_{\alpha}}\hat{n}_{\mathbf{r}}+e^{-i\mathbf{q\cdot e}_{\alpha}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}\right)],$ (8)
@@
$H_{\text{FB}}\left|2,q\right\rangle=H_{\text{FB}}\left|\psi_{2}\right\rangle+2\cos q\left(\sum_{l}^{N}\hat{b}_{l}^{{\dagger}}\hat{b}_{l}\hat{b}_{l+1}^{{\dagger}}\hat{b}_{l+1}\right)\left|2,q\right\rangle.$
@@
$\left\langle\psi_{n}\right|\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\left|\psi_{n}\right\rangle=e^{-i\mathbf{q\cdot R}}\frac{\left(N-n\right)n}{N(N-1)},$ (21)
@@ 4 Resonant scattering of non-Hermitian impurity
$p_{j}(t)=\left|\frac{\hat{a}_{j}\left|\phi(t)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2}=\left|\frac{\hat{a}_{j}e^{-iH_{\text{HB}}t}\left|\phi(0)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2},$ (39)
@@
$P(t)=\frac{1}{n}\sum_{j}p_{j}(t)=\left|\frac{e^{-iH_{\text{HB}}t}\left|\phi(0)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2},$ (40)
respectively. We plot the two quantities in Fig. 1(a) and (b). The parameters of the initial wavepackets and the driven system are given in the captions. The profiles of eveloved states and the total probabilities indicate the perfect absorption for the resonant incident wavepackets with $q=q_{\text{c}}$ for both one- and two-boson cases, in accord with the previous theoretical analysis. The results indicate that the time evolutions for the cases with $n=1$ and $n=2$ share the similar profiles. There are two implications: (i) The resonant absorptions demonstrate the dynamic signature of the EP. (ii) The similarity of the dynamic behavior shows that the NN interaction indeed plays a role in canceling out the scattering effect arising from the hardcore interaction.
Table 1: The structures of energy levels for $10$-site open chain with different $q$ and filling particle number $n$. We list the numbers of coalescing states $n_{\text{CS}}$, of order $n_{\text{OR}}$, and the numbers of complex levels $n_{\text{CM}}$ in the form ($n_{\text{CM}}$, $n_{\text{OR}}\times n_{\text{CS }}$). It indicates that all the energy levels are real and all the systems contain a single $2$ order coalescing state.
@@
Table 2: The same as Table 1 but for $N_{1}\times N_{2}=5\times 3$ lattice with different ($q_{1}$, $q_{2}$). We take open boundary condition in $q_{1}$-direction and periodic boundary condition in $q_{2}$-direction. It indicates that all the systems contain complex levels and multiple $2$ order coalescing states.
@@
Figure 2: (a1)-(a4) present the time evolution of fidelity of $n$-particle one-dimensional system with $q=m\pi/N$ where m ranges from $1$ to $4$. (b1)-(b4) present the time evolution of probability of the same system. The slope of dashed line as reference is $2$ and it is nearly parallel to other colored lines. Here we only plot the results with a single 2 order coalescing state. The initial state is $\prod_{j=1}^{n}\hat{a}_{j}^{\dagger}\left|0\right\rangle$ and the size of system $N=10$, where the number n is indicated in the panel.
@@ 5 Dynamic generation of condensate states
Condensate state as macroscopic quantum state has significance both in theoretical and experimental physics. The stability of such states is crucial in practice. In general, it can be prepared as ground state by discreasing the temperature. Dynamical generation of nonequilibrium steady condensate states is another way, which has been received much attention recently. The advantage of an EP system is that a coalescing state is also a target state for a long-time evolution of various initial states. In the above, we have shown that the state $\left|\psi_{n}\right\rangle$ is also a coalescing state of the Hamiltonian $H(q_{1}^{\mathrm{c}})$. However, the order of EP is unkown, maybe depends on the particle numbers $n$, and the numbers of dimensions $n_{\mathrm{d}}$ at which the resonant boundary is taken.
@@ A Condensate eigenstates with ODLRO
]$\ H\left|\psi_{n}\right\rangle=n\sum\limits_{\alpha=1}^{3}V_{\alpha}\left|\psi_{n}\right\rangle.$ (A.7)
@@
$\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\left|\psi_{n}\right\rangle=$ $\frac{1}{\Omega_{n}}\{nA^{n-1}[e^{-i\mathbf{q\cdot r}}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\hat{a}_{\mathbf{r}}^{\dagger}$
$+e^{-i\mathbf{q\cdot}(\mathbf{r}+\mathbf{R})}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\hat{a}_{\mathbf{r}+\mathbf{R}}^{\dagger}]\}\left|0\right\rangle$
@@ B Coalescing condensate states
$H_{\text{HB}}^{1}=$ $\sum_{j=1}^{N-1}\left(\hat{a}_{j}^{\dagger}\hat{a}_{j+1}+\text{H.c.}+\cos q\hat{n}_{j}\hat{n}_{j+1}\right)$
$+e^{-iq}\hat{a}_{1}^{\dagger}\hat{a}_{1}+e^{iq}\hat{a}_{N}^{\dagger}\hat{a}_{N},$ (B.4)

### v3

@@ 1 Introduction
On the other hand, recent years have seen a growing interest in non-Hermitian descriptions of condensed-matter systems [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. It has been shown that the interplay between non-Hermiticity and interaction can give rise to exotic quantum many-body effect, ranging from non-Hermitian extensions of Kondo effect [15, 27], many-body localization [21], Fermi surface in coordinate space [28], to fermionic superfluidity [18, 29]. The cooperation between the non-Hermiticity and interaction may lead to rich quantum phases due to the peculiarity of non-Hermitian system.
@@ 2 Model and condensate states
$H=\sum\limits_{\alpha=1}^{3}J_{\alpha}\sum_{\mathbf{r}}\frac{1}{2}\hat{a}_{\mathbf{\ r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{e}_{\alpha}}+\mathrm{H.c.}+\sum\limits_{\alpha=1}^{3}V_{\alpha}\sum_{\mathbf{r}}\hat{n}_{\mathbf{r}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}+\sum\limits_{\alpha=1}^{3}\sum_{\mathbf{r}}\left(\mu_{\alpha}\hat{n}_{\mathbf{r}}\delta_{1,m_{\alpha}}+\mu_{\alpha}^{\ast}\hat{n}_{\mathbf{r}}\delta_{N_{\alpha},m_{\alpha}}\right),$ (1)
@@
$h_{\mathbf{r}}^{\alpha}=J_{\alpha}\left[\frac{1}{2}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{e}_{\alpha}}+\mathrm{H.c.}+\cos(\mathbf{q\cdot e}_{\alpha})(\hat{n}_{\mathbf{r}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}-\hat{n}_{\mathbf{r}}-\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}})+\frac{1}{2}\left(e^{i\mathbf{q\cdot e}_{\alpha}}\hat{n}_{\mathbf{r}}+e^{-i\mathbf{q\cdot e}_{\alpha}}\hat{n}_{\mathbf{r}+\mathbf{e}_{\alpha}}\right)\right],$ (8)
@@
$H_{\text{FB}}\left|2,q\right\rangle=H_{\text{FB}}\left|\psi_{2}\right\rangle+2\cos q\left(\sum_{l}^{N}\hat{b}_{l}^{{\dagger}}\hat{b}_{l}\hat{b}_{l+1}^{{\dagger}}\hat{b}_{l+1}\right)\left|2,q\right\rangle.$ (18)
@@
$\left\langle\psi_{n}\right|\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\left|\psi_{n}\right\rangle=e^{-iq\cdot\mathbf{R}}\frac{\left(N-n\right)n}{N(N-1)},$ (22)
@@ 4 Resonant scattering of non-Hermitian impurity
$p_{j}(t)=\left|\frac{\hat{a}_{j}\left|\phi(t)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2}=\left|\frac{\hat{a}_{j}e^{-iH_{\text{HB}}^{\infty}t}\left|\phi(0)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2},$ (40)
@@
$P(t)=\frac{1}{n}\sum_{j}p_{j}(t)=\left|\frac{e^{-iH_{\text{HB}}^{\infty}t}\left|\phi(0)\right\rangle}{\left|\phi(0)\right\rangle}\right|^{2},$ (41)
respectively.
We plot the two quantities $p_{j}(t)$ and $P(t)$ in Fig. 1(a) and (b) for selected systems with two types of initial states. The parameters of the initial wavepackets and the driven systems are given in the captions of the figure. We consider the driven Hamiltonians on a finite site chain with the parameter $q_{\text{c}}=\pi/10$. The initial states are taken as: (a) a single-boson wavepacket and (b) a two-boson wavepacket, respectively. The center momenta of the wavepackets are taken as $q=q_{\text{c}}$, $q_{\text{c}}/2$, and $2q_{\text{c}}$, respectively. According to the analysis above, both the single- and two-boson incident wavepackets with the resonant momentum $q_{\text{c}}$ should be approximately absorbed by the ending non-Hermitian impurity. From Fig. 1(a1)-(a3), one can observe that there is no reflection pattern in (a1) for the resonant wavepacket. In contrast, evident interference fringes are observed near the end of the chain in cases (a2) and (a3), where the incident wavepackets are off-resonant. Such patterns result from the interference between incident and reflected wavepackets. From Fig. 1(a4), one can see that the probabilities for each case reduce significantly when the wavepackets reach the impurity, exhibiting absorption behavior. Notably, the absorption is minimal when the initial wavepacket carries the resonant momentum $q_{\text{c}}$. This result is consistent with the patterns observed in Fig. 1(a1)-(a3). Actually, the plots in Fig. 1(a1)-(a4) are provided to compare the results with those for two-boson incident wavepackets, which are shown in Fig. 1(b1)-(b4). We observe that the two sets of patterns have the similar profiles. The results indicate perfect absorption for resonant incident wavepackets with $q=q_{\text{c}}$ in both the one- and two-boson cases, which is in accordance with the previous theoretical analysis. Then, on the one hand, the results indicate that the underlying mechanism for the coalescing condensate states in an open chain is the reflectionless absorption of many-particle wavepackets with resonant momentum by the non-Hermitian boundary. These resonant absorptions demonstrate the dynamic signature of the EP. On the other hand, it reveals an interesting dynamic behavior of a many-particle wavepacket: it behaves as a single-particle wavepacket. The NN resonant interaction indeed plays a role in canceling out the scattering effect arising from the hardcore interaction.
Table 1: The structures of energy levels for $10$-site open chain with different $q$ and filling particle number $n$. We list the numbers of coalescing states $n_{\text{CS}}$, of order $n_{\text{OR}}$, and the numbers of complex levels $n_{\text{CM}}$ in the form ($n_{\text{CM}}$, $n_{\text{OR}}\times n_{\text{CS }}$). It indicates that all the energy levels are real and all the systems contain $2$-order coalescing states.
@@
Table 2: The same as Table 1 but for $N_{1}\times N_{2}=5\times 3$ lattice with different ($q_{1}$, $q_{2}$). We take open boundary condition in $q_{1}$-direction and periodic boundary condition in $q_{2}$-direction. It indicates that all the systems contain complex levels and multiple $2$-order coalescing states.
@@
Figure 2: (a1)-(a4) present the time evolution of fidelity of $n$-particle one-dimensional system with $q=m\pi/N$ where m ranges from $1$ to $4$. (b1)-(b4) present the time evolution of probability of the same system. The slope of dashed line as reference is $2$ and it is nearly parallel to other colored lines. Here we only plot the results with a single 2 order coalescing state. The initial state is $\prod_{j=1}^{n}\hat{a}_{j}^{\dagger}\left|0\right\rangle$ and the size of system $N=10$, where the number n is indicated in the panel. The results show that the unentangled initial states can indeed evolve to the corresponding target with high fidelity.
@@ 5 Dynamic generation of condensate states
Condensate state as macroscopic quantum state has significance both in theoretical and experimental physics. The stability of such states is crucial in practice. In general, it can be prepared as ground state by decreasing the temperature. Dynamical generation of nonequilibrium steady condensate states is another way, which has been received much attention recently. The advantage of an EP system is that a coalescing state is also a target state for a long-time evolution of various initial states. In the above, we have shown that the state $\left|\psi_{n}\right\rangle$ is also a coalescing state of the Hamiltonian $H(q_{1}^{\mathrm{c}})$. However, the order of EP is unknown, maybe depends on the particle numbers $n$, and the numbers of dimensions $n_{\mathrm{d}}$ at which the resonant boundary is taken.
@@
which is the measure of the distance between the final state and the condensate states $\left|\psi_{n}(q)\right\rangle$. We plot the fidelity $F(t)$ in Fig. 2 as a function of $t$ for selected systems and many-particle initial states. We consider the driven Hamiltonians on $10$-site chains with the system parameters $q=m\pi/10$, where $m$ ranges from $1$ to $4$. For $m=1$ and $3$, we consider the sectors with $n=2$, $3$, $4$, and $5$. For $m=2$ and $4$, we only consider the sectors with $n=2$ and $3$. According to the results listed in Tables 1, the corresponding energy levels are all real and the systems contain only one $2$-order EP, with the coalescing condensate states in the form
$\left|\psi_{n}\right\rangle=\frac{1}{\Omega_{n}}\left(\sum_{j=1}^{10}\hat{a}_{j}^{\dagger}e^{-im\pi j/10}\right)^{n}\left|0\right\rangle.$ (48)
The $n$-boson initial states are taken in the explicit form
$\left|\phi\left(0\right)\right\rangle=\hat{a}_{1}^{\dagger}\left|0\right\rangle,\hat{a}_{1}^{\dagger}\hat{a}_{2}^{\dagger}\left|0\right\rangle,\hat{a}_{1}^{\dagger}\hat{a}_{2}^{\dagger}\hat{a}_{3}^{\dagger}\left|0\right\rangle,\hat{a}_{1}^{\dagger}\hat{a}_{2}^{\dagger}\hat{a}_{3}^{\dagger}\hat{a}_{4}^{\dagger}\left|0\right\rangle,\hat{a}_{1}^{\dagger}\hat{a}_{2}^{\dagger}\hat{a}_{3}^{\dagger}\hat{a}_{4}^{\dagger}\hat{a}_{5}^{\dagger}\left|0\right\rangle,$ (49)
@@ A Condensate eigenstates with ODLRO
$H\left|\psi_{n}\right\rangle=n\sum\limits_{\alpha=1}^{3}V_{\alpha}\left|\psi_{n}\right\rangle.$ (A.7)
@@
$\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\left|\psi_{n}\right\rangle=$ $\frac{1}{\Omega_{n}}\left\{nA^{n-1}[e^{-i\mathbf{q\cdot r}}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\hat{a}_{\mathbf{r}}^{\dagger}+e^{-i\mathbf{q\cdot}(\mathbf{r}+\mathbf{R})}\hat{a}_{\mathbf{r}}^{\dagger}\hat{a}_{\mathbf{r}+\mathbf{R}}\hat{a}_{\mathbf{r}+\mathbf{R}}^{\dagger}]\right\}\left|0\right\rangle$
@@ B Coalescing condensate states
$H_{\text{HB}}^{1}=$ $\sum_{j=1}^{N-1}\left(\hat{a}_{j}^{\dagger}\hat{a}_{j+1}+\text{H.c.}+\cos q\hat{n}_{j}\hat{n}_{j+1}\right)+e^{-iq}\hat{a}_{1}^{\dagger}\hat{a}_{1}+e^{iq}\hat{a}_{N}^{\dagger}\hat{a}_{N},$ (B.4)